OpenStudy (anonymous):

give a number a, try to calculate square root of a by going with f(x)= x^2-a, and using newton's method to find an approximation of the number x* that makes f(x*)=0. what the update formula is ,

OpenStudy (anonymous):

@TuringTest can you help me on this question?

OpenStudy (asnaseer):

What do you already know about Newtons method?

OpenStudy (anonymous):

\[x_{n+1}=x_{n}-f(x)/f'(x)\]

OpenStudy (anonymous):

\[f(x)\neq 0\]

OpenStudy (asnaseer):

that is correct - so where are you stuck?

OpenStudy (anonymous):

i don't understand what is question mean? how can i express this?

OpenStudy (asnaseer):

ok, first of all, it is asking you to select some number (represented by 'a')

OpenStudy (anonymous):

i got the \[x* = \sqrt{a}\]

OpenStudy (asnaseer):

so first pick some number

OpenStudy (asnaseer):

it is asking you to use newton's method to find the square root of a number that you select

OpenStudy (asnaseer):

so you need to first select some number

OpenStudy (anonymous):

ok. i pick a = 10

OpenStudy (asnaseer):

good, so now we represent this in f(x) as:\[f(x)=x^2-10\]

OpenStudy (asnaseer):

and we need to find a value for x such that f(x) = 0 as this will give us the square root of 10 - agreed?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

that is what i got

OpenStudy (asnaseer):

ok, so first we /guess/ at a solution what is your first guess at the solution?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

it's half value of [0,10]

OpenStudy (asnaseer):

good - so we can say \(x_0=5\) and then proceed to iterate to the answer (within some defined accuracy) using newton's formula:\[x_{n+1}=x_{n}-f(x_n)/f'(x_n)\]now notice that you need f'(x) in this formula, so first work out f'(x) given:\[f(x)=x^2-10\]

OpenStudy (anonymous):

f'(x)=2x

OpenStudy (asnaseer):

great! so if we now substitute f(x) and f'(x) into newton's formula, we get:\[x_{n+1}=x_{n}-\frac{x_n^2-10}{2x_n}=x_n-\frac{x_n^2}{2x_n}+\frac{10}{2x_n}=x_n-0.5x_n+\frac{5}{x_n}=0.5x_n+\frac{5}{x_n}\]make sense?

OpenStudy (anonymous):

yes! i got the point! thank you ! i have one more question. would you mind help me?

OpenStudy (asnaseer):

unfortunately I need to go now - but just post your new question in the list to the left and I'm sure someone will come to your aid :)

OpenStudy (anonymous):

thank you!

OpenStudy (asnaseer):

yw :)